Be sure to answer all parts.

A solution contains 0.253 mol of NaCl dissolved in 8.86 mol of water.

(a) What is the mole fraction of NaCl?


(b) What is the molality of the solution?
Answer #1

(a) If a solution contains x mole of 'a' component , y mole of 'b' component ,then mole fraction a component in the mixture is expressed as a/(a+b) . Similarly for b component b/(a+b).

So, mole fraction is an expreesion for concentration but it is unit less because of it is a ratio of two same quantity.

Mole fraction if NaCl is 0.253/(0.253+8.86) = 0.0278 .

(b) Molality of a solution is defined as the number of moles of solute present per 1Kg ( or 1000 g) of solvent molecule.

We know that 1 mole H2O = 18 g.

So, 8.86 mole water = (8.86*18) gm = 159.48 g.

159.48 g water cotains 0.253 mole of NaCl.

1000 g water contains (0.253/159.48)*1000 mole = 1.59

Molality of the solution is 1.59 .